Solve the system by the substitution method.
x + 5y = 5
3x - 8y = -8

Answer by rapaljer(4551)

First, solve for one of the variables in one of the equations. The easiest is to solve for x in the first equation:
x= 5-5y

Substitute this back into the second equation in place of the x:
3x - 8y = -8
3(____) - 8y = -8
3(5-5y) - 8y = -8
15 - 15y - 8y = -8
15-23y = -8
15-15 -23y = -8-15
-23y = -23
y=1

Now, substitute y = 1 back into the EASIEST equation to find x, which would be the x= 5-5y equation
x=5-5y
x=5 -5*1
x=0

Check in the OTHER equation (not the one you just used to find x!):
3x-8y=-8
3(0)-8(1)=-8
-8=-8 It checks!

R^2 at SCC